Whats the kinetic energy of a car that's mass is m=1000kg that is traveling at a speed of 100km per hour


1
Expert's answer
2021-09-29T09:55:07-0400

By definition, the kinetic energy is:


"K = \dfrac{mv^2}{2}"

where "v = 100km/h = 100/3.6\space m/s" is the speed of the car, and "m = 1000kg" is its mass. Thus, obtain:


"K = \dfrac{1000\cdot (100/3.6)^2}{2} \approx 3.86\times 10^5J"

Answer. "3.86\times 10^5J".